Two particles each of mass m are moving in horizontal circle with same angular speed . If both string are of same length then ratio of tension in string T1/T2

Grade:11

1 Answers

PRASANNA
6 years ago
Tension is provided by centripetal force : T = mw2r
T2 = mw2(2r) ------------ (1)
WKT : T1 – T2 = mw2r
from (1)
T1 – 2mw2r = mw2r
T1 = 3mw2r ------------------ (2)
 
from (1) and (2)
T1/T2 = 3/2